How many moles of gaseous arsine (AsH3) occupy 0.0400 L at STP? What is the density of gaseous arsine?
1
Expert's answer
2019-12-03T08:25:11-0500

"PV=nRT"

Where; pressure P=101325pa

            Volume V=o.0400L=4*10^-5m3

            Temperature T=273.15K

            R= 8.314 J/K

            Number of moles =n

"n =(PV)\/(RT)"

n = (101325*4*10^-5)/ (8.314*273.15).

n =1.7857*10^-3moles

1mole of AsH3 has a molar mass of 77.95g.

1 mole =77.95g ,1.7857*10^-3moles will have a mass ,m of 0.139.

density, "d=m\/v"

d =0.139*10^-3kg/4*10^-5m3

d =3.475kg/m3.